Two discs with identical moments of inertia (initial inertia) begin separated one on the other rotating around the same axis. The top one is rotating in favor of the clock hands at 2.0 rad / s, and the second one is rotating against the clock hands at 6.0 rad / s. If we release the one above and the two identical discs collide inelastically sticking. What will be the angular velocity just after the collision when they are joined?

- la -06
Answer #1

By conservation of angular momentum,

final angular momentum = initial angular momentum

If*Wf = I*w1 + I*w2 where I is moment of inertia of each disc

2I* Wf = I*2 + I*-6

2Wf = 2-6 = -4

Wf = -4/2 = -2 rad/s

  angular velocity will be 2 rad/s against the clock